Grade 11Trigonometry

. If the point (3, -2) is transformed to (-2, 1) when the origin is shifted to (x,y) then x-y is equal to
9

Jagruthi
5 Years agoGrade 11

1 Answer

Priyansu Samal
5 Years ago
If A+B+C=180°
Then, prove tanA+tanB+tanC=tanA.tanB.tanC
Ana-         L.H.S.
          A+B+C=180° ;A+B =180°-C
         tan(A+B)=tan(180°-C)
         tanA+tanB/1-tanA.tanB=-tanC
         tanA.tanB.tanC (RHS)